A gas has a volume of 3.28 L at a temperature of 28.80 oC and pressure of 0.37 atm. What is its volume if the pressure changes to 1.67 atm and the temperature changes to 43.81 oC?


A gas has a volume of 14.20 L at a temperature of 372.18 K and pressure of 3.54 atm. What is its temperature if the pressure changes to 1.25 atm and the volume changes to 5.60 L?


A gas is compressed to a volume of 16.80 L at a pressure of 2.70 atm and a temperature of 290.23 K. If the initial pressure was 1.39 atm and the initial volume was 31.21 L, what was the initial temperature?


What was the initial pressure on a gas if its volume changes from 4.15 L to 11.23 L, its temperature changes from 13.70 oC to 45.00 oC and its final pressure is 0.70 atm?


What was the initial volume of a gas if its pressure changes from 3.49 atm to 1.73 atm, its temperature changes from 140.67 K to 203.05 K, and its final volume is 8.78 L?



1
Expert's answer
2021-05-18T06:51:02-0400

Answer :-

for all parts we use

"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"


1)"V=\frac{0.37\times3.28\times316.81}{308.18\times 1.67}=0.74L"

we changed temperature to kelvin by adding 273


2)"T=\frac{1.25\times5.6\times372.18}{14.2\times3.54}=51.82K"


3) T="\frac{290.23\times1.39\times 31.21}{16.80\times2.70}=277.57K"


4)"P=\frac{286.7\times11.23\times0.70}{4.15\times 318}=1.72atm"

we changed temperature to kelvin by adding 273


5)"V=\frac{140.67\times8.78\times 1.73}{3.39\times 203.05 }=3.10L"